# RPO: Retrieval Preference Optimization for Robust Retrieval-Augmented Generation

### v1

@@ Abstract
While Retrieval-Augmented Generation (RAG) has exhibited promise in utilizing external knowledge, its generation process heavily depends on the quality and accuracy of the retrieved context. Large language models (LLMs) struggle to evaluate the correctness of non-parametric knowledge retrieved externally when it differs from internal memorization, leading to knowledge conflicts during response generation. To this end, we introduce the Retrieval Preference Optimization (RPO), a lightweight and effective alignment method to adaptively leverage multi-source knowledge based on retrieval relevance. An implicit representation of retrieval relevance is derived and incorporated into the reward model to integrate retrieval evaluation and response generation into a single model, solving the problem that previous methods necessitate the additional procedure to assess the retrieval quality. Notably, RPO is the only RAG-dedicated alignment approach that quantifies the awareness of retrieval relevance in training, overcoming mathematical obstacles. Experiments on four datasets demonstrate that RPO outperforms RAG by 4-10% in accuracy without any extra component, exhibiting its robust generalization.
@@
Shi-Qi Yan, Zhen-Hua Ling National Engineering Research Center of Speech and Language Information Processing, University of Science and Technology of China, Hefei, China sqyan01@mail.ustc.edu.cn, zhling@ustc.edu.cn
@@ 3.1 RAG Formulation
$\begin{cases}y_{n+p}=\pi(q,R)\\ R=D^{r},\end{cases}$ (1)
where $y_{n+p}$ means the answer for the question $q$ that has access to the retrieved results, i.e., all retrieved context $D^{r}$. LLMs autonomously select either parametric or non-parametric knowledge for response generation.
@@ 3.2 Knowledge Conflict
$\begin{cases}y_{p}=\pi(q,R)\\ R=\phi,\end{cases}$ (2)
where $y_{p}$ means the answer without any retrieved context, i.e. null set in the equation above, representing the response with parametric knowledge for $q$. Note that if the parametric knowledge and retrieved non-parametric knowledge are different, i.e., knowledge conflict arises, the generator in RAG should make a decision on which knowledge to be referred to. If the knowledge from the retrieved context is adopted, the answer would be vary from $y_{p}$. Based on this situation, we filtered the non-parametric answers $y_{n}$ from $y_{n+p}$ Ultimately, we can detect knowledge conflict and filter non-parametric answers by:
@@ 4 Why DPO is Limited to Apply to RAG
Considering the integrated retrieved context in the input when applied to RAG, the ultimate optimization objective of RLHF and DPO can be formulated as :
@@ 5 Methodology
Figure 2: An overview of RPO at training. In phase 1, given a question and the retrieved documents, two answers $(y_{p},y_{n})$ are generated by the frozen language model $\pi$. After comparing with the golden answers that are annotated, instances that involve knowledge conflict are filtered for supervised fine-tuning. In phase two, the tine-tuned LLM is prompted to generate a pair of answers again, and the instances with knowledge conflict are filtered as the training set of RPO.
@@ 5.1 Theoretically Analysis
$\begin{split}r(x,y,R)=&\beta\log{\frac{\pi(y\mid x,R)}{\pi_{\text{ref}}(y\mid x,R)}}\\ &+\beta\log{\frac{\pi(R\mid x)}{\pi_{\text{ref}}(R\mid x)}}+\beta\log{Y(x)},\\ \end{split}$ (8)
@@
$\begin{split}&\mathcal{L}_{\text{RPO{}}}=-\mathbb{E}\bigg{[}\log\sigma\bigg{(}\underbrace{\beta\log{\frac{\pi_{\theta}(y_{w}|x,D^{r})}{\pi_{\text{ref}}(y_{w}|x,D^{r})}}}_{\text{(a)preferred generation reward}}\\ &\underbrace{-\beta\log{\frac{\pi_{\theta}(y_{l}|x,D^{r})}{\pi_{\text{ref}}(y_{l}|x,D^{r})}}}_{\text{(b)dispreferred generation reward}}\underbrace{{\pm\frac{\beta}{\left|D^{r}\right|}\log{\frac{\pi_{\theta}(D^{r}\mid x)}{\pi_{\text{ref}}(D^{r}\mid x)}}}}_{\text{(c)retreival reward}}\bigg{)}\bigg{]},\end{split}\vspace{-1mm}$ (9)
@@
Dataset($\mathbf{D}$) : $\mathcal{X}$ (Input Questions), $\mathcal{Y}$ (Output Labels), =$\mathbb{C}=\{D_{1},D_{2},...,D_{N}\}$ (Documents)
@@ 2      $y_{p}=\pi(x)$
6     $\mathbf{D}_{\text{SFT}}=$Conflict_Collection($\mathbf{D}$, Condition:$\text{Acc}(y_{n+p})+\text{Acc}(y_{p})=1$ )
@@ 9           $y_{p}=\pi_{\text{SFT}}(x)$
13          $\mathbf{D}_{\text{RPO{}}}=$Conflict_Collection($\mathbf{D}$, Condition:$\text{Acc}(y_{n+p})+\text{Acc}(y_{p})=1$)
@@ 5.2 Training Overview
We first construct the preference pairs that are adopted for supervised fine-tuning (SFT) and RPO, aimed at enhancing the model’s awareness to leverage retrieved non-parametric knowledge adaptively. Given an instance from the dataset $(x,y)\in\mathbf{D}$, we respectively instruct the model to generate responses with and without retrieval $(y_{n+p},y_{p})$ as illustrated in Section 3.2. Two subsets sampled from $\mathbf{D}$ are constructed to collect preference pairs. In the first subset $\mathbf{D}^{1}$, our goal is to continually enhance the model’s ability to read and comprehend the retrieved context. Instances are sampled where the model fails to answer the questions directly, while correctly generating the responses with retrieval, i.e., $\text{Acc}(y_{n+p})>\text{Acc}(y_{p})$. To further confirm that $y_{n+p}$ refers to the retrieved knowledge, i.e. $y_{n+p}=y_{n}$, we solely select samples where the ground truths are contained in the retrieved context. The second subset $\mathbf{D}^{2}$ focuses on mitigating the over-reliance of the model on the retrieved knowledge. We select the instances where the model could have responded correctly while being affected by the retrieved knowledge and generating incorrect answers, i.e., $\text{Acc}(y_{p})>\text{Acc}(y_{n+p})$. Note that interference due to incorrectness is caused by the introduced non-parametric knowledge, $y_{n+p}$ can be approximately regarded as a non-parametric answer $y_{n}$. It helps the model to reconsider whether and when to utilize the non-parametric knowledge before generation. Ultimately, combine both subsets and obtain the training set, $\mathbf{D}_{train}=\mathbf{D}^{1}\cup\mathbf{D}^{2}$, which consists of samples that involve knowledge conflict.
@@ 6 Experiments
Table 1: Overall evaluation results on the test sets of four datasets. Results are separated based on the generation LLMs. The Column Adaptive Category indicates the category of the method if it belongs to adaptive RAG. # API/LM Calls means the number of times that an API or an LM is called during an inference. Bold numbers indicate the best performance among all methods and LLMs. * indicates the results that are directly cited from the papers, otherwise results are reproduced by us with the consistent retrieval results.
@@
We conducted experiments to extensively demonstrate RPO’s advancement and adaptability to RAG-based approaches and its generalizability across various tasks.
@@ 6.2 Baselines
LLaMA2-7B approaches utilized the vanilla or instruction-tuned LLaMA2-7B model for response generation. Self-RAG (Asai et al., 2023) that tuned the LLaMA2 on the instruction-tuning data containing several sets of reflection tokens which were labeled by GPT-4 (OpenAI, 2023).
@@
Others means that LLMs except LLaMA2 and LLaMA3 are adopted as the generator. Specifically, AstuteRAG was reproduced in this experiment on ChatGPT, which iteratively filtered and revised the knowledge before generation.
@@ 6.3 Results
Table 1 presents the results on four datasets. We briefly mark the categories of the listed adaptive RAG baselines in the table. To showcase the efficiency of our RPO in computational overhead during the inference phase, the estimated API call or LLM inference times are presented as well. From these results, we can conclude the following findings:
First, the proposed method significantly outperformed previous baselines that involve adaptive retrieval, reaching state-of-the-art. Specifically, as shown in table 1, RPO outperformed RAG by margins of 7.4% accuracy on PopQA, 10.6% accuracy on NQ, 8.6% accuracy on TriviaQA, and 3.7% accuracy on RGB when based on LLaMA3-8B-instruct, as well as by margins of 7.0% accuracy on PopQA, 23.3% accuracy on NQ, 5.1% accuracy on TriviaQA, and 5.7% on RGB when based on LLaMA2-hf-7b. Compared with the currently advanced adaptive RAG methods, RPO has generally outperformed in all the benchmarks. The advancements in our method greatly illustrate the effectiveness of preference optimization, showing the significance of overcoming the knowledge conflict.
@@ 6.4 Ablation Study
Table 2: Ablation study for removing retrieval-awareness, preference optimization, and SFT phases respectively on the PopQA dataset in terms of accuracy.
@@
Since two phases including supervised fine-tuning and preference optimization have adopted during the training stage, and both of them help to enhance retrieval awareness and overcome knowledge conflict. We perform ablation studies to confirm the contribution of each stage in our RPO. The fine-tuning and preference optimization phases are removed specifically in the experiment and the results are evaluated on the benchmarks. Results in Table 2 demonstrate that the performance dropped when removing either phases, revealing the significance.
@@ 6.6 Knowledge Selection Performance
In this section, we compare RPO with previous training strategies in terms of knowledge selection performance. Since the original purpose in this paper is to address the knowledge conflict, activate the capability of evaluating the retrieval, and align with the more accurate answer. We do further analysis on the issue of knowledge conflict before and after RPO. The results in figure 3 reveal a consistent advancement in all clusters to evaluate the knowledge and select the correct autonomously. Especially in Cluster B and C, which involve knowledge conflicts, RPO has shown a significant improvement in knowledge selection.

### v2

@@ Abstract
While Retrieval-Augmented Generation (RAG) has exhibited promise in utilizing external knowledge, its generation process heavily depends on the quality and accuracy of the retrieved context. Large language models (LLMs) struggle to evaluate the correctness of non-parametric knowledge retrieved externally when it differs from internal memorization, leading to knowledge conflicts during response generation. To this end, we introduce the Retrieval Preference Optimization (RPO), a lightweight and effective alignment method to adaptively leverage multi-source knowledge based on retrieval relevance. An implicit representation of retrieval relevance is derived and incorporated into the reward model to integrate retrieval evaluation and response generation into a single model, solving the problem that previous methods necessitate the additional procedure to assess the retrieval quality. Notably, RPO is a RAG-dedicated alignment approach that quantifies the awareness of retrieval relevance in training, first overcoming mathematical obstacles. Experiments on four datasets demonstrate that RPO outperforms RAG by 4-10% in accuracy without any extra component, exhibiting its robust generalization.
@@
Shi-Qi Yan1, Quan Liu2and Zhen-Hua Ling1 1National Engineering Research Center of Speech and Language Information Processing, University of Science and Technology of China 2State Key Laboratory of Cognitive Intelligence, iFLYTEK Research sqyan01@mail.ustc.edu.cn, quanliu@iflytek.com, zhling@ustc.edu.cn
@@ 3.1 RAG Formulation
$y_{n+p}=\pi(x,R)|_{R=D^{r}},\vskip-2.84526pt$ (1)
where $y_{n+p}$ means the answer for the question $x$ that has access to the retrieved results, i.e., all retrieved context $D^{r}$. LLMs autonomously select either parametric or non-parametric knowledge for response generation.
@@ 3.2 Knowledge Conflict
$y_{p}=\pi(x,R)|_{R=\phi},$ (2)
where $y_{p}$ means the answer without any retrieved context, i.e. null set in the equation above, representing the response with parametric knowledge for $x$. Note that if the parametric knowledge and retrieved non-parametric knowledge are different, i.e., knowledge conflict arises, the generator in RAG should make a decision on which knowledge to be referred to. If the knowledge from the retrieved context is adopted, the answer would be vary from $y_{p}$. Based on this situation, we filtered the non-parametric answers $y_{n}$ from $y_{n+p}$ Ultimately, we can detect knowledge conflict and filter non-parametric answers by:
@@ 4 Why DPO is Limited to Apply to RAG
Considering the integrated retrieved context in the input when applied to RAG, the ultimate optimization objective of PPO-based methods such as RLHF and DPO can be formulated as :
@@ 5 Methodology
Figure 2: An overview of RPO at training. In phase 1, given a question and the retrieved documents, two answers $(y_{p},y_{n})$ are generated by the frozen language model $\pi$. After comparing with the golden answers, instances that involve knowledge conflict are filtered for supervised fine-tuning. In phase two, the fine-tuned LLM is prompted to generate a pair of answers again, and the instances with knowledge conflict are filtered as the training set of RPO.
@@ 5.1 Theoretically Analysis
[…]
@@
$\begin{split}&\mathcal{L}_{\text{RPO{}}}=-\mathbb{E}\bigg[\log\sigma\bigg(\underbrace{\beta\log{\frac{\pi_{\theta}(y_{w}|x,D^{r})}{\pi_{\text{ref}}(y_{w}|x,D^{r})}}}_{\text{(a)preferred generation reward}}\\ &\underbrace{-\beta\log{\frac{\pi_{\theta}(y_{l}|x,D^{r})}{\pi_{\text{ref}}(y_{l}|x,D^{r})}}}_{\text{(b)dispreferred generation reward}}\underbrace{{\pm\frac{\beta}{\left|D^{r}\right|}\log{\frac{\pi_{\theta}(D^{r}\mid x)}{\pi_{\text{ref}}(D^{r}\mid x)}}}}_{\text{(c)retreival reward}}\bigg)\bigg],\end{split}$ (9)
@@
Dataset($\mathbf{D}$) : $\mathcal{X}$ (Input Questions), $\mathcal{Y}$ (Output Labels), $\mathbb{C}=\{D_{1},D_{2},...,D_{N}\}$ (Documents)
@@ 2   $y_{p}=\pi(x)$
[…]
@@ 9     $y_{p}=\pi_{\text{SFT}}(x)$
[…]
@@ 5.2 Training Overview
We first construct the preference pairs adopted for supervised fine-tuning (SFT) and RPO, aimed at enhancing the model’s awareness to leverage retrieved non-parametric knowledge adaptively. Given an instance from the dataset $(x,y)\in\mathbf{D}$, we respectively instruct the model to generate responses with and without retrieval $(y_{n+p},y_{p})$ as illustrated in Section 3.2. Two subsets sampled from $\mathbf{D}$ are constructed to collect preference pairs. In the first subset $\mathbf{D}^{1}$, our goal is to continually enhance the model’s ability to read and comprehend the retrieved context. Instances are sampled where the model fails to answer the questions directly, while correctly generating the responses with retrieval, i.e., $\text{Acc}(y_{n+p})>\text{Acc}(y_{p})$. To further confirm that $y_{n+p}$ refers to the retrieved knowledge, i.e. $y_{n+p}=y_{n}$, we solely select samples where the ground truths are contained in the retrieved context. The second subset $\mathbf{D}^{2}$ focuses on mitigating the over-reliance of the model on the retrieved knowledge. We select the instances where the model could have responded correctly while being affected by the retrieved knowledge and generating incorrect answers, i.e., $\text{Acc}(y_{p})>\text{Acc}(y_{n+p})$. Note that interference due to incorrectness is caused by the introduced non-parametric knowledge, $y_{n+p}$ can be approximately regarded as a non-parametric answer $y_{n}$. It helps the model to reconsider whether to utilize the non-parametric knowledge before generation. Ultimately, combine both subsets and obtain the training set, $\mathbf{D}_{train}=\mathbf{D}^{1}\cup\mathbf{D}^{2}$, which consists of samples that involve knowledge conflict.
@@ 6 Experiments
Table 1: Overall evaluation results on the test sets of four datasets. Results are separated based on the generation LLMs. The Column Adaptive Category indicates the category of the method if it belongs to adaptive RAG. # API/LM Calls means the number of times that an API or an LM is called during an inference. Bold numbers indicate the best performance among all methods and LLMs. †indicates that due to the cost, only a part of the test set is evaluated. * indicates the results that are directly cited from the papers, otherwise results are reproduced by us with the consistent retrieval results.
@@
We conducted experiments to extensively demonstrate RPO’s advancement and adaptability to RAG-based approaches and their generalizability across various tasks.
@@ 6.2 Baselines
LLaMA2-7B approaches utilized the vanilla or instruction-tuned LLaMA2-7B model for response generation. (1) RAG + SFT directly tuned the model with the instances that involve knowledge conflict. (2) RAG + DPO tuned the model with SFT in phase 1, while tuning the model with DPO rather than RPO in Phase 2. Conflict collection is implemented in both SFT and DPO before training to ensure comparability. (3) Self-RAG (Asai et al., 2023) that tuned the LLaMA2 on the instruction-tuning data containing several sets of reflection tokens which were labeled by GPT-4 (OpenAI, 2023), while (4) CRAG (Yan et al., 2024) that evaluated the quality of the retrieval and selectively corrected the retrieved context with the web search.
@@
Commercial APIs refers to the approaches that import commercial LLMs for text generation. We introduce the methods driven by commercial APIs for reference to benchmark the broader effectiveness and efficiency of our proposed RPO. Specifically, AstuteRAG (Wang et al., 2024) was reproduced in this experiment on ChatGPT, which iteratively filtered and revised the knowledge before generation.
@@ 6.3 Results
Table 1 presents the results on four datasets. We briefly mark the categories of the listed adaptive RAG methods in the table. To showcase the efficiency of RPO in computational overhead during the inference phase, the estimated API call or LLM inference times are presented as well. From these results, we can conclude the following findings:
First, the proposed method significantly outperformed previous baselines that involve adaptive retrieval, reaching state-of-the-art. Specifically, as shown in table 1, RPO outperformed RAG by margins of 6.4% accuracy on PopQA, 10.6% accuracy on NQ, 8.6% accuracy on TriviaQA, and 3.7% accuracy on RGB when based on LLaMA3-8B-instruct, as well as by margins of 7.0% accuracy on PopQA, 23.3% accuracy on NQ, 5.1% accuracy on TriviaQA, and 5.7% on RGB when based on LLaMA2-hf-7b. Compared with the currently advanced adaptive RAG methods, RPO has generally outperformed in all the benchmarks. The advancements in our method greatly illustrate the effectiveness of preference optimization, showing the significance of overcoming the knowledge conflict.
@@ 6.4 Ablation Study
Table 2: Ablation study for removing retrieval-awareness, preference optimization, and SFT phases respectively on the PopQA dataset in terms of accuracy. w̃/o RR means that the retrieval reward term is removed for optimization, while w̃/o PO means that the model is trained without preference optimization.
@@
Given that our training pipeline incorporates two distinct phases—supervised fine-tuning and preference optimization and both of which contribute to enhancing retrieval awareness and mitigating knowledge conflict, we conduct ablation studies to evaluate the individual contribution of each phase within our RPO framework. The fine-tuning and preference optimization phases are removed specifically in the experiment and the results are evaluated on the benchmarks. It is worth noting that, since the retrieval reward term in Equ. (9) is the biggest difference between DPO and RPO, RPO without the retrieval reward term (RR) can be equivalent to a DPO model. Similarly, RPO without preference optimization represents models trained solely via supervised fine-tuning, omitting the subsequent alignment stage. Results in Table 2 demonstrate that the performance dropped when removing either phases, revealing the significance.
6.5 Robustness to Low-Quality Retrieval
Table 3: The robustness of each training strategy to low-quality retrieval in the PopQA dataset, where all retrieval information is incorrect.
Acc in Low-Quality Retrieval
LLaMA2-7B-hf
RAG            18.6 (0.0%)
SFT            19.5 (+4.8%)
DPO            19.3 (+3.7%)
RPO            23.5 (+26.3%)
As illustrated above, one of the primary objectives in this paper is to improve the ability of LLMs to select accurate information amidst knowledge conflicts. It frequently occurs in a low-quality retrieval environment, posing significant challenges for prior methods. Therefore, to further evaluate the robustness of RPO to low-quality retrieval, we simulate this environment by assessing the performance of LLMs when provided only with incorrect information. Results in Table 3 reveal the performance degradation of various methods under the condition of erroneous retrieval context in PopQA. Although all methods inevitably suffer from performance degradation, our RPO still maintains a superior performance. The experiments further demonstrate that unlike DPO, which exhibits limitations and potential biases when applied to RAG, RPO can effectively evaluate the correctness of the retrieved context during the response generation.
@@ 6.7 Knowledge Selection Performance
In this section, we compare RPO with previous training strategies in terms of knowledge selection performance. Further analysis is conducted on the issue of knowledge conflict before and after RPO. The results in figure 3 reveal a consistent advancement in all clusters to evaluate the knowledge and select the correct autonomously. Besides, we found that the ability of the LLM to select knowledge can be even worse after SFT. In Cluster B and C, which involve knowledge conflicts, SFT does not achieve a positive advancement, while RPO has shown a significant improvement in knowledge selection.
@@ 7 Conclusion
Limitations
While we primarily proposed to improve the RAG framework with a dedicated alignment method, whether a better reward function exists requires further study. Although we make an effort to prevent reward hacking during the experiments, the intended objective can still not be fully fulfilled. In addition, since the model is only trained on NQ, the training data could not cover various domains, leading to potential bias. Future work will further explore a more flexible and robust rewarding strategy for RAG.
@@ A.3 Derivation of RPO’s Reward Model
Appendix B Experiment Details
B.1 Details of the Datasets
RPO was evaluated on four datasets, which are in public domain and licensed for research purposes, including:
PopQA Mallen et al. (2023) is a short-form generation task. Generally, only one entity of factual knowledge is expected to be answered for each single question. In our experiments, we exactly followed the setting in the previous work Asai et al. (2023) which evaluated methods on a long-tail subset consisting of 1,399 rare entity queries whose monthly Wikipedia page views are less than 100.
Natural Questions (NQKwiatkowski et al. (2019) is a benchmark for question answering research that contains real user questions issued to Google search, and answers found from Wikipedia by annotators. Annotations include long answers (usually a paragraph of text) and short answers (one or more entities), which are marked as null if there is no answer on the page. Additionally, NQ contains 307,372 training examples, 7,830 examples for development, and we withold a further 7,842 examples for testing. Only short answers are adopted in our experiments.
TriviaQA Joshi et al. (2017) is a reading comprehension dataset containing over 650K question-answer-evidence triples. TriviaQA includes 95K question-answer pairs authored by trivia enthusiasts and independently gathered evidence documents, six per question on average, that provide high quality distant supervision for answering the questions.
Retrival-Augmented Generation Benchmark (RGBChen et al. (2024) is a benchmark that chooses to aggregate the latest news. Different basic abilities of LLMs are evaluated according to the common challenges in RAG, including noise robustness, negative rejection, information integration and counterfactual robustness.
B.2 Experimental Setup
We use the package vllm for inference, and the parameter settings are listed below:
temperature=0.0
top_p=1.0
max_tokens=100
skip_special_tokens=false
The model was trained on 4*A100 in our experiment, and the SFT was implemented with the hyperparameter settings below:
n_epochs=1
batch_size=4
gradient_accumulation_steps=32
mixed_precision=bf16
max_seq_length=2048
warmup_ratio=0.03
learning_rate=2e-5
weight_decay=0.0,
while RPO strictly followed the hyperparameters used in Rafailov et al. (2023).